Orange trees ~ In the 1960s, a botanist was conducting a research on orange trees. He wanted to figure out the relationship between the growth of trees and their ages. The botanist gathered a random sample of 35 orange trees and recorded the circumference of the tree in mm and the age of the tree in days.
Circumference of tree is the X variable and Age of tree is the Y variable in this scenario.

Using the sample data, the botanist finds the equation of the estimated regression line for predicting the Y variable to be yˆ =116.9093 + 7.8636 x

A tree called Amber has a circumference of 108.6834 mm.

What is the estimated/predicted Age for the tree called Amber? Give your answer to 4 decimal places.

Note: Numbers are randomized for each instance of this question. Use the numbers given above.

Your Answer:

Answer
Answer #1

Given that ,

yˆ =116.9093 + 7.8636 x

We plug x = 108.6834 we get ,

yˆ =116.9093 + 7.8636 * 108.6834 = 971.5521

estimated/predicted Age = 971.5521